OpenStudy (anonymous):

The distance between the end of a lamppost’s shadow and the top of the lamppost is 22 feet. The angle measure of the shadow is 15°. How tall is the lamppost to the nearest tenth of a foot?

OpenStudy (anonymous):

haha very easy simply treat it as a triangle... using pythagoras theorem we can solve it..

OpenStudy (anonymous):

 sin theta = o/h sin15 = x/22 22sin15=x x=5.7ft

OpenStudy (turingtest):

so we can use the sine function to solve for the opposite side.

OpenStudy (anonymous):

oww i got 5.9